OpenStudy (shadowcain):

Indicate which property is illustrated in Step 3 Step 1 5 + (12 + 5) + 0 = 5 + (5 + 12) + 0 Step 2 = (5 + 5) + (12 + 0) Step 3 = (5 + 5) + 12 Step 4 = 10 + 12 A. distributive property B. identity property of addition C. associative property of addition D. identity property of multiplication

OpenStudy (freckles):

which operation is happening? multiplication or addition?

OpenStudy (shadowcain):

C. associative property of addition

OpenStudy (freckles):

we have 12+0=12 12 is the same as 12 does that sound like an identity or associative property?

OpenStudy (freckles):

and I meant to say 12+0 is the same as 12* associative property we are changing where the grouping symbols are

OpenStudy (shadowcain):

The identity property for addition tells us that zero added to any number is the number itself. Zero is called the "additive identity." The identity property for multiplication tells us that the number 1 multiplied times any number gives the number itself.

OpenStudy (freckles):

and we have 12+0=12... so you think the answer is...

OpenStudy (freckles):

hint: that operation definitely isn't multiplication

OpenStudy (shadowcain):

I know it's not multiplication. You are explaining step 2 (5 + 5) + (12 + 0) The question is about Step 3

OpenStudy (freckles):

yes in step 3 they have wrote 12+0 as 12

OpenStudy (freckles):

you are almost there you even defined this property above

OpenStudy (shadowcain):

B. identity property of addition

OpenStudy (freckles):

yes!!! The identity property for addition tells us that zero added to any number is the number itself. Zero is called the "additive identity." This is equivalent to saying a+0=a The identity property for multiplication tells us that the number 1 multiplied times any number gives the number itself. a*1=a

OpenStudy (shadowcain):

thanks freckles!!